# JointMotion: Joint Self-supervision for Joint Motion Prediction

### v1

@@ Abstract
We present JointMotion, a self-supervised learning method for joint motion prediction in autonomous driving. Our method includes a scene-level objective connecting motion and environments, and an instance-level objective to refine learned representations. Our evaluations show that these objectives are complementary and outperform recent contrastive and autoencoding methods as pre-training for joint motion prediction. Furthermore, JointMotion adapts to all common types of environment representations used for motion prediction (i.e., agent-centric, scene-centric, and pairwise relative), and enables effective transfer learning between the Waymo Open Motion and the Argoverse 2 Forecasting datasets. Notably, our method improves the joint final displacement error of Wayformer, Scene Transformer, and HPTR by 3%, 7%, and 11%, respectively.
Keywords:
Self-supervised learning representation learning multi-modal pre-training motion prediction
Figure 1: JointMotion. (a) Connecting motion and environments: Our scene-level objective learns joint scene representations by matching embeddings of motion sequences ($Z^{M}$) and environment context ($Z^{E}$). (b) Masked instance modeling: Our instance-level objective includes a complete set of perception predictions: agent position ($x$, $y$), dimensions ($l$, $w$, $h$), acceleration (a), velocity (v), yaw angle ($\theta_{\text{yaw}}$), temporal order ($n_{\text{t}}$), and class ($c_{\text{a}}$); lane node position (x, y) and class ($c_{\text{l}}$); traffic light position ($x,y$), temporal order ($n_{\text{t}}$), and class ($c_{\text{tl}}$).
@@ 1 Introduction
Self-supervised learning (SSL), where supervision signals are generated from unlabeled data, by design excels in applications with large amounts of unlabeled data and limited labeled data (e.g., [4, 8, 15]). Recently, SSL methods combined with supervised fine-tuning outperform plain supervised learning with the same amount of data [5, 16] and shorter training times [12]. This makes SSL a versatile choice to improve existing methods in a wide range of applications.
In this work, we focus on improving joint motion prediction for autonomous driving. Motion prediction aims to predict the future motion of traffic agents given their past motion and environment context, such as lane networks and traffic light states. Therefore, motion prediction is essential for estimating future interaction amongst traffic agents [6] and subsequently assessing the risk of planned trajectories [23]. The majority of recent motion prediction methods perform marginal prediction, where future motion is predicted for each agent individually (e.g., [26, 31, 21]). Joint motion prediction predicts scene-wide motion modes, i.e. a mode captures a joint set of future motion sequences for multiple agents within a traffic scene, enabling interaction modeling.
Since joint motion prediction requires scene-wide representations, we propose a scene-level SSL objective connecting motion and environment. We complement this by an instance-level SSL objective that refines these representations, enhancing overall prediction accuracy. As scene-level objective, we match all past motion sequences in a scene to the corresponding environment context, such as map data and traffic light states. Thereby, a model learns which sets of motion sequences are likely in a given environment, including interaction amongst agents. As instance-level objective, we reconstruct masked motion sequences, masked lane polylines, and masked sequences of traffic light states. Unlike related work [7, 9], our instance-level objective includes a complete set of object features (agent dimensions, position, orientation, speed and acceleration profiles, and temporal order) to match to the environment context, rather than just positions. Hence, our main contributions are the two jointly optimized SSL objectives that accelerate the training and improve the precision of motion prediction models:
  1. 1.
    Connecting motion and environments via similarity learning, see Fig. 1 (a),
  2. 2.
    Masked instance modeling with a complete set of object features, see Fig. 1 (b).
@@ 2 Related work
SSL for motion prediction in autonomous driving. Inspired by the success of SSL in computer vision (e.g., [8, 15, 16]) and natural language processing (e.g., [4, 24]), related methods apply SSL to pre-train motion prediction models for autonomous driving. PreTraM [33] learns a joint embedding space for trajectory and map data via contrastive learning. During pre-training, the similarity of trajectory and map embeddings from the same traffic scene is maximized, while the similarity to embeddings from other scenes is minimized (i.e., negative examples). This objective is misleading when similar scenes are sampled in a mini-batch, e.g., the map embeddings of two four-way stops should be similar, but are erroneously used as negative examples against each other. The concurrent works Traj-MAE [7], Forecast-MAE [9], and RMP [34] are masked autoencoding methods for pre-training motion prediction models. They mask trajectory and/or lane tokens and learn to reconstruct them. Traj-MAE and Forecast-MAE represent motion sequences as trajectory polylines, where each polyline point represents only the agents’ position omitting available features such as velocity and acceleration profiles, agent dimensions, yaw angles, and temporal order (see Fig. 1 (b)). Furthermore, they do not include traffic light states as environment context. RMP limits the scope of pre-training to motion sequences and omits environment context, such as lane data and traffic light states, entirely. SSL-Lanes [3] extends lane masking with the objectives of distance to intersection prediction, maneuver and success/failure classification. Although these objectives are well adapted to motion prediction, they require non-trivial heuristics (e.g., for clustering maneuvers).
@@
Joint motion prediction for autonomous driving. Joint motion predictions aims to predict the joint distribution of future motion sequences over multiple agents in a traffic scene. Thus, a predicted motion mode represents a scene-wide set of motion sequences with one sequence per agent. Scene Transformer [22] is an encoder-decoder transformer model for joint motion prediction. The encoder learns global scene-centric representations (i.e., agent, lane, and traffic light features), the decoder transforms these embeddings into joint motion modes. Global position embeddings are used for the learned scene-centric representations, which are more difficult to learn than pairwise relative position embeddings (cf. [36, 10]). MotionLM [25] reformulates motion prediction as language prediction task and learns a vocabulary of discrete motion vectors. Joint prediction modes are generated by jointly decoding sequences of motion vectors for multiple agents. MotionDiffuser [18] performs joint motion prediction as conditional denoising diffusion process. A scene-wide set of noisy trajectories (i.e., positions disturbed by Gaussian noise) is transformed by a denoiser into a scene-wide set of trajectories that approximates the ground truth joint prediction mode. The learned denoiser is conditioned on environment context such as lane data and traffic light states. At inference, trajectories are sampled from pure noise. Both MotionLM and MotionDiffuser use an agent-centric encoder [21], which leads to repeated computation since a traffic scene is encoded for each agent individually. To achieve a scene-wide context for a shared decoder, these methods concatenate multiple agent-centric embeddings with positional and rotation information into a common reference frame. Furthermore, their best performing variants exhibit high inference latency and are not real-time capable (MotionLM:  250 ms [^1] This latency is linearly interpolated, as inference latency is only reported up to the second best/largest model (see Table 8 in the Appendix of [25]), MotionDiffuser: 409 ms).
@@
Marginal motion prediction with auxiliary interactive prediction objectives. A marginal motion mode represents a single motion sequence for one agent. Many recent methods extend marginal prediction models with modules for interactive prediction objectives (i.e., dense [14, 27] or conditional prediction [29, 20]). In such methods, the prediction modules are still trained in a marginal manner, where training targets are marginal motion modes per agent. The additional modules are trained to combine marginal predictions by minimizing overlap between them. Thus, these methods do not learn to model scene-wide joint motion modes from ground up, but how to combine marginal modes with less overlap. This limits interaction modeling to the learned re-combination of marginal motion predictions.
@@ 3 Method
We present our method for self-supervised pre-training of motion prediction models in two steps. First, we describe our scene-level SSL objective (connecting motion and environments), then, we describe our agent-level SSL objective (masked instance modeling).
@@ 3.1 Connecting motion and environments
Joint motion prediction requires scene-wide representations to decode joint modes, which include predictions of multiple agents within a scene. Therefore, we propose pre-training motion prediction models by connecting motion and environments. This scene-level objective aims to learn a joint embedding space for motion sequences and environment context (see Fig. 1 (a)). Thus, a model implicitly learns which motion is likely in a given environment, including traffic rules and interaction amongst agents.
@@
In detail, we use the past motion of all agents within a scene to generate a scene-level motion embedding ($Z^{M}$ in Fig. 1) and combine lane data and traffic light states to a corresponding environment embedding ($Z^{E}$ in Fig. 1). These embeddings are generated by modality-specific encoders (i.e., for motion, lanes, and traffic lights) followed by global average pooling and an MLP-based projector (Proj in Fig. 1). We perform average pooling on the intermediate embeddings ($H^{M}$ and $H^{E}$ in Fig. 1) to be invariant to variations in the number of agents and the complexity of environments. Following [13], we use two separate MLPs with LayerNorm and ReLU activations as projectors, each having a hidden dimension of 2048 and an output dimension of 256. We use the modality-specific encoders of recent motion prediction models (e.g., [22, 36]) without modifications and remove the additional projector after pre-training. The joint embedding space is learned by similarity learning via redundancy reduction. Following [35, 30], we reduce the redundancy of vector elements per embedding (i.e., for $Z^{M}$ and $Z^{E}$ individually) and maximize their similarity by approximating the cross-correlation matrix of $Z^{M}$ and $Z^{E}$ to the corresponding identity matrix. The redundancy reduction term ensures that individual embedding elements capture different features. Therefore, it prevents model collapse with trivial solutions for all embeddings across all scenes (e.g., zero vectors). Non-trivial yet identical solutions are not explicitly prevented, but are unlikely, as empirically shown in [35]. In contrast to [33], our scene-level objective does not require negative examples, which are difficult to define in this context. Unlike [30], this objective maximizes the similarity of embeddings from different modalities (i.e., motion and environment) rather than augmented views of the same modality, removing the requirement to develop suitable augmentations.
3.2 Masked instance modeling
@@
Scene-level representations are well suited to provide an overview of traffic scenes, but lack details at instance level. For example, the exact position of traffic agents in the past or lane curvatures. Therefore, we combine our scene-level objective with the instance-level objective of masked instance modeling to refine learned representations.
@@
Inspired by masked sequence modeling [4, 16], we mask elements of past motion sequences, lane polylines, and past traffic light states and learn to reconstruct them from non-masked elements and environment context. As shown in Fig. 1, we represent traffic agents with 10 features rather than just past positions (cf. [7, 9]). In detail, we reconstruct agent positions ($x$, $y$), dimensions ($l$, $w$, $h$), acceleration (a), velocity (v), yaw angle ($\theta_{\text{yaw}}$), temporal order ($n_{\text{t}}$), and classes ($c_{\text{A}}$, i.e., vehicles, cyclists, and pedestrians). For lanes, we reconstruct positions ($x,y$) and lane classes ($c_{\text{L}}$). For traffic light state sequences, we reconstruct positions ($x$, $y$), temporal order ($n_{\text{t}}$), and state classes ($c_{\text{TL}}$, i.e., green, yellow, and red). Following [36], positions, dimensions, accelerations, velocities, and yaw angles are represented as float values and temporal order and agent classes as boolean one-hot encodings.
@@
For models that use late fusion mechanisms (e.g., [22, 36]), we use the modality-specific encoders to generate embeddings per modality (see Fig. 2 (a)). Afterwards, we concatenate these embeddings and use a shared local decoder to reconstruct masked sequence elements from non-masked elements and context from other modalities. As local decoder, we use transformer blocks with PreNorm, local attention [2] with an attention window of 32 tokens, 8 attention heads, rotary positional embeddings [28], and feed forward layers with an input and hidden dimension of 256 and 1024. For models that employ early fusion mechanisms (e.g., [21]), we use learned queries and a shared decoder to reconstruct the input sequences (see Fig. 2 (b)). Such models learn a compressed latent representation for multi-modal input. Therefore, we use learned queries in the same number as input tokens to decompress these representations and reconstruct the input sequences. We use a regular cross-attention mechanism between the learned queries and compressed latent representations and a local self-attention mechanism within the set of learned queries. The resulting transformer blocks have the same hyperparameters as in the late fusion setup. For both variants, we use random attention masks for masking and a masking ratio 60%. As training target, we minimize the Huber loss between the reconstructed sequences ($S^{R}$) and the input sequences ($S^{I}$): $\mathcal{L}_{\text{recon}}=\lambda_{\text{A}}\mathcal{L}_{\text{A}}(S^{R}_{\text{A}},S^{I}_{\text{A}})+\lambda_{\text{L}}\mathcal{L}_{\text{L}}(S^{R}_{\text{L}},S^{I}_{\text{L}})+\lambda_{\text{TL}}\mathcal{L}_{\text{TL}}(S^{R}_{\text{TL}},S^{I}_{\text{TL}})$. If not specified otherwise, we set $\lambda_{\text{A}}=\lambda_{\text{L}}=\lambda_{\text{TL}}=1$.
@@ 4 Experiments
4.1 Datasets and evaluation metrics
The Waymo Open Motion Dataset (WOMD) [11] is comprised of over 1.1 million data points extracted from 103,000 urban or suburban driving scenarios, spanning 20 seconds each. The state of object-agents includes attributes like position, dimensions, velocity, acceleration, orientation, angular velocity, and the status of turn signals. Each data point captures 1 second of past followed by 8 seconds of future data. We resample this time interval with 10Hz.
Evaluation metrics. Following [11], we use the mean average precision (mAP), the average displacement error (minADE), and the final displacement error (minFDE) to evaluate motion predictions. All metrics are computed using the minimum mode for $K=6$ modes. Accordingly, the metrics for the mode closest to the ground truth are measured. We use the official challenge website to compute metrics on WOMD. For the AV2 dataset, we use the evaluation software provided by [36]. Therefore, joint modes (best scene-wide mode) are evaluated on the interactive splits and marginal modes (best mode for each agent individually) on the regular splits.

### v2

@@ Abstract
We present JointMotion, a self-supervised pre-training method for joint motion prediction in self-driving vehicles. Our method jointly optimizes a scene-level objective connecting motion and environments, and an instance-level objective to refine learned representations. Scene-level representations are learned via non-contrastive similarity learning of past motion sequences and environment context. At the instance level, we use masked autoencoding to refine multimodal polyline representations. We complement this with an adaptive pre-training decoder that enables JointMotion to generalize across different environment representations, fusion mechanisms, and dataset characteristics. Notably, our method reduces the joint final displacement error of Wayformer, HPTR, and Scene Transformer models by 3%, 8%, and 12%, respectively; and enables transfer learning between the Waymo Open Motion and the Argoverse 2 Motion Forecasting datasets. Code: https://github.com/kit-mrt/future-motion
Keywords: Self-supervised learning, representation learning, multimodal pre-training, motion prediction, data-efficient learning
Figure 1: JointMotion. (a) Connecting motion and environments: Our scene-level objective learns joint scene representations via non-contrastive similarity learning of motion sequences $M$ and environment context $E$. (b) Masked polyline modeling: Our instance-level objective refines learned representations via masked autoencoding of multimodal polyline embeddings (i.e., motion, lane, and traffic light data).
@@ 1 Introduction
Self-supervised learning (SSL) [1] by design excels in applications with large amounts of unlabeled data and limited labeled data (e.g., [2, 3, 4]). However, recent SSL methods combined with supervised fine-tuning outperform plain supervised learning with the same amount of data [5, 6] and shorter training times [7]. This makes SSL a versatile choice to improve existing methods in a wide range of applications.
In this work, we focus on improving joint motion prediction for self-driving vehicles. Motion prediction aims to predict the future motion of traffic agents given their past motion and environment context, such as lane data and traffic light states. Therefore, motion prediction is essential for estimating future interaction among traffic agents [8] and subsequently assessing the risk of planned trajectories [9, 10]. The majority of recent motion prediction methods perform marginal prediction, where future motion is predicted for each agent individually (e.g., [11, 12]). In contrast, joint motion prediction predicts scene-wide motion modes (i.e., capturing a joint set of future motion sequences for multiple agents), which enables interaction modeling.
Since joint motion prediction requires scene-wide representations, we propose a scene-level SSL objective connecting motion and environment. We combine this with an instance-level SSL objective that refines learned representations, enhancing overall prediction accuracy. As scene-level objective, we match past motion sequences in a scene to the corresponding environment context, such as map data and traffic light states. Thereby, a model learns which sets of motion sequences are likely in a given environment, including interaction among agents. As instance-level objective, we reconstruct masked polylines representing past motion sequences, lanes, and past traffic light states. We complement this with an adaptive pre-training decoder, which leads to the high generalizability of our method.
Overall, our main contributions are the design choices that enable our method to generalize across
  1. (a)
    environment representations (i.e., scene-centric, pairwise relative, and agent-centric) enabled by our complementary scene- and instance-level pre-training objectives,
  2. (b)
    information fusion mechanisms (i.e., early and late fusion) enabled by the combination of our adaptive pre-training decoder and our instance-level objective, and
  3. (c)
    dataset characteristics (i.e., varying sequence lengths and feature sets) enabled by the transferability of learned representations.
@@ 2 Related work
SSL for motion prediction in self-driving vehicles. Inspired by the success of SSL in computer vision (e.g., [3, 4, 6]) and natural language processing (e.g., [2, 13, 14]), related methods apply SSL to pre-train motion prediction models for self-driving. PreTraM [15] learns a joint embedding space for trajectory and map data via contrastive learning. During pre-training, the similarity of trajectory and map embeddings from the same traffic scene is maximized, while the similarity to embeddings from other scenes is minimized (i.e., negative examples). This objective is counterproductive when similar scenes are sampled in a mini-batch, e.g., the map embeddings of two four-way stops should be similar, but are erroneously used as negative examples against each other. The concurrent works Traj-MAE [16], Forecast-MAE [17], and RMP [18] are masked autoencoding methods for pre-training motion prediction models. They mask trajectory and/or lane tokens and learn to reconstruct them. Traj-MAE and Forecast-MAE represent motion sequences as trajectory polylines, where each polyline point represents only the agents’ position omitting available features such as velocity and acceleration profiles, agent dimensions, yaw angles, and temporal order (see Fig. 1 (b)). Furthermore, they do not include traffic light states as environment context. SSL-Lanes [19] extends lane masking with the objectives of distance to intersection prediction, maneuver and success/failure classification. Although these objectives are well adapted to motion prediction, they require non-trivial heuristics (e.g., for clustering maneuvers).
@@
Joint motion prediction for self-driving vehicles. Joint motion prediction aims to predict the joint distribution of future motion sequences over multiple agents in a traffic scene. Thus, a predicted motion mode represents a scene-wide set of motion sequences with one sequence per agent. Scene Transformer [20] is an encoder-decoder transformer model for joint motion prediction. The encoder learns global scene-centric representations (i.e., agent, lane, and traffic light features), the decoder transforms these embeddings into joint motion modes. Global position embeddings are used for the learned scene-centric representations, which are more difficult to learn than pairwise relative position embeddings (cf. [21, 22]). MotionLM [23] reformulates motion prediction as language prediction task and learns a vocabulary of discrete motion vectors. Joint prediction modes are generated by autoregressively decoding sequences of motion vectors for multiple agents. MotionDiffuser [24] performs joint motion prediction as conditional denoising diffusion process. A set of noisy trajectories (i.e., positions disturbed by Gaussian noise) is transformed by a denoiser into a set of trajectories that approximates the ground truth joint prediction mode. The learned denoiser is conditioned on environment context such as lane data and traffic light states. Both MotionLM and MotionDiffuser use an agent-centric encoder [12], which leads to repeated computation since a traffic scene is encoded for each agent individually. Furthermore, their best performing variants exhibit high inference latency (MotionLM:  250 ms [^1] This latency is linearly extrapolated, as inference latency is only reported up to the second best/largest model (see Table 8 in the Appendix of [23]), MotionDiffuser: 409 ms).
@@
Marginal motion prediction with auxiliary interaction prediction objectives. A marginal motion mode represents a single motion sequence for one agent. Many recent methods extend marginal prediction models with modules for interaction prediction objectives (i.e., dense [25, 26, 27] or conditional prediction [28, 29]). In such methods, the main prediction modules are still trained in a marginal manner, where training targets are marginal motion modes per agent. The additional modules are trained to combine marginal predictions or the corresponding latent representations by minimizing overlap between them. Thus, these methods do not learn to model scene-wide joint motion modes from ground up, but how to combine marginal modes with less overlap. This limits interaction modeling to the learned re-combination of marginal motion predictions.
@@ 3 Method
We present our method for self-supervised pre-training of motion prediction models in two steps. First, we describe our scene-level SSL objective (connecting motion and environments), then, our instance-level SSL objective (masked polyline modeling) and the adaptive pre-training decoder.
@@ 3.1 Connecting motion and environments
Joint motion prediction requires scene-wide representations to decode joint modes, which include predictions of multiple agents within a scene. Therefore, we propose pre-training motion prediction models by connecting motion and environments (CME). This scene-level objective aims to learn a joint embedding space for motion sequences and environment context (see Figure 1 (a)). Thus, a model implicitly learns which motion is likely in a given environment, including traffic rules and interaction among agents.
@@
In detail, we use the past motion of all agents within a scene to generate a scene-level motion embedding ($Z^{M}$ in Figure 1) and combine lane data and traffic light states to a corresponding environment embedding ($Z^{E}$ in Figure 1). These embeddings are generated by modality-specific encoders (i.e., for motion, lanes, and traffic lights) followed by global average pooling and an MLP-based projector (Proj in Figure 1). We perform average pooling on the intermediate embeddings ($H^{M}$ and $H^{E}$ in Figure 1) to be invariant to variations in the number of agents and the complexity of environments. Following [30], we use two separate MLPs with LayerNorm and ReLU activations as projectors, each having a hidden dimension of 2048 and an output dimension of 256. We use the modality-specific encoders of recent motion prediction models (e.g., [20, 21]) without modifications and remove the additional projector after pre-training. The joint embedding space is learned by similarity learning via redundancy reduction. Following [31, 32], we reduce the redundancy of vector elements per embedding (i.e., for $Z^{M}$ and $Z^{E}$ individually) and maximize their similarity by approximating the cross-correlation matrix $C$ of $Z^{M}$ and $Z^{E}$ to the corresponding identity matrix:
$\mathcal{L}_{\text{CME}}=\lambda_{\text{red}}\sum_{i}\sum_{j\neq i}{C_{ij}}^{2}+\sum_{i}(1-C_{ii})^{2},$
where $i,j$ index the vector dimension of the embeddings $Z$. The redundancy reduction term is scaled by $\lambda_{\text{red}}$ and ensures that individual embedding elements capture different features. Therefore, it prevents representation collapse with trivial solutions for all embeddings across all scenes (e.g., zero vectors). Non-trivial yet identical solutions are not explicitly prevented, but are unlikely, as empirically shown in [31]. In contrast to [15], our scene-level objective does not require negative examples, which are difficult to define in this context. Unlike [32], this objective maximizes the similarity of embeddings from different modalities (i.e., motion and environment) rather than augmented views of the same modality, removing the requirement to develop suitable augmentations.
3.2 Masked polyline modeling with adaptive decoding
@@
Scene-level representations are well suited to provide an overview of traffic scenes, but lack instance-level details. For example, the exact position of traffic agents in the past or lane curvatures. Therefore, we combine our scene-level objective with the instance-level objective of masked polyline modeling (MPM) to refine learned representations.
@@
Inspired by masked sequence modeling [2, 6], we mask elements of polylines representing past motion sequences, lanes, and past traffic light states and learn to reconstruct them from non-masked elements and environment context. As shown in Figure 1, we represent traffic agents with 10 features rather than just past positions (cf. [16, 17]). In detail, we reconstruct agent positions ($x$, $y$), dimensions ($l$, $w$, $h$), acceleration a, velocity v, yaw angle $\theta_{\text{yaw}}$, temporal order $n_{\text{t}}$, and classes $c_{\text{A}}$ (i.e., vehicles, cyclists, and pedestrians). For lanes, we reconstruct positions ($x,y$) and lane classes $c_{\text{L}}$. For traffic light state sequences, we reconstruct positions ($x$, $y$), temporal order $n_{\text{t}}$, and state classes $c_{\text{TL}}$ (i.e., green, yellow, and red). Following [21], positions, dimensions, accelerations, velocities, and yaw angles are represented as float values and temporal order and agent classes as boolean one-hot encodings.
@@
For models with late or hierarchical fusion mechanisms (e.g., [20, 21]), we use the modality-specific encoders to generate embeddings per modality (see Figure 2 (a)). Afterwards, we concatenate these embeddings and use a shared local decoder to reconstruct masked sequence elements from non-masked elements and context from other modalities. As local decoder, we use transformer blocks with PreNorm [33], local attention [34] with an attention window of 32 tokens, 8 attention heads, rotary positional embeddings [35], and FeedForward layers with an input and hidden dimension of 256 and 1024.
For models that employ early fusion mechanisms (e.g., [12]), we use learned queries and a shared decoder to reconstruct the input sequences (see Figure 2 (b)). Such models learn a compressed latent representation for multi-modal input. Therefore, we use learned queries in the same number as input tokens to decompress these representations and reconstruct the input sequences. We use a regular cross-attention mechanism between the learned queries and compressed latent representations and a local self-attention mechanism within the set of learned queries. The resulting transformer blocks have the same hyperparameters as in the late fusion setup.
@@
For both variants, we use random attention masks for masking and a masking ratio 60%. As training target, we minimize the Huber loss between the reconstructed polylines and the input polylines $\mathcal{L}_{\text{MPM}}=\lambda_{\text{A}}\mathcal{L}_{\text{A}}+\lambda_{\text{L}}\mathcal{L}_{\text{L}}+\lambda_{\text{TL}}\mathcal{L}_{\text{TL}}$. If not specified otherwise, we set $\lambda_{\text{A}}=\lambda_{\text{L}}=\lambda_{\text{TL}}=1$.